# DICE: End-to-end Deformation Capture of Hand-Face Interactions from a Single Image

### v1

@@ Abstract
Reconstructing 3D hand-face interactions with deformations from a single image is a challenging yet crucial task with broad applications in AR, VR, and gaming. The challenges stem from self-occlusions during single-view hand-face interactions, diverse spatial relationships between hands and face, complex deformations, and the ambiguity of the single-view setting. The first and only method for hand-face interaction recovery, Decaf [97], introduces a global fitting optimization guided by contact and deformation estimation networks trained on studio-collected data with 3D annotations. However, Decaf suffers from a time-consuming optimization process and limited generalization capability due to its reliance on 3D annotations of hand-face interaction data. To address these issues, we present DICE, the first end-to-end method for Deformation-aware hand-face Interaction reCovEry from a single image. DICE estimates the poses of hands and faces, contacts, and deformations simultaneously using a Transformer-based architecture. It features disentangling the regression of local deformation fields and global mesh vertex locations into two network branches, enhancing deformation and contact estimation for precise and robust hand-face mesh recovery. To improve generalizability, we propose a weakly-supervised training approach that augments the training set using in-the-wild images without 3D ground-truth annotations, employing the depths of 2D keypoints estimated by off-the-shelf models and adversarial priors of poses for supervision. Our experiments demonstrate that DICE achieves state-of-the-art performance on a standard benchmark and in-the-wild data in terms of accuracy and physical plausibility. Additionally, our method operates at an interactive rate (20 fps) on an Nvidia 4090 GPU, whereas Decaf requires more than 15 seconds for a single image. Our code will be publicly available upon publication.
@@ 1 Introduction
Hand-face interaction is a common behavior observed up to 800 times per day across all ages and genders [98]. Therefore, faithfully recovering hand-face interactions with plausible deformations is an important task given its wide applications in AR/VR [87, 36, 113], character animation [88, 136], and human behavior analysis [66, 29, 74]. Given the speed requirement of downstream applications like AR/VR, fast and accurate 3D reconstruction of hand-face interactions is highly desirable. However, several challenges make monocular hand-face deformation and interaction recovery a cumbersome task: 1) the self-occlusion involved in hand-face interaction, 2) the diversity of hand and face poses, contacts, and deformations, and 3) ambiguity in the single-view setting. Most existing methods [91, 75] only reconstruct the hand [92] and face [57] meshes, or unified as a whole body [67, 82] with the body parts, without capturing contacts and deformations. A seminal advance, Decaf [97], recovers hand-face interactions with deformations and contacts taken into account. However, it requires time-consuming optimization, which takes more than 15 seconds per image, rendering it unsuitable for interactive applications. The iterative fitting process of Decaf relies on an accurate estimation of hand and face keypoints and contacts on the hand and face surfaces, which could fail when significant occlusion is present in the image; See Fig. 8 in the Appendix D.1. Additionally, Decaf cannot scale up their training to fruitful hand-face interaction data in the wild, as they require 3D ground-truth annotations, i.e., contact labels and deformations.
@@
Figure 1: Our method, DICE, is the first end-to-end approach that captures hand-face interaction and deformation from a monocular image. (a) Decaf validation dataset. (b) In-the-wild images. (c) Use-cases in VR.
@@
To tackle the issues above, we present DICE, the first end-to-end approach for Deformation-aware hand-face Interaction reCovEry from a monocular image. We use a Transformer-based model with the attention mechanism to effectively capture hand-face relationships. Motivated by the global nature of the pose and shape of hand and face and the local nature of the deformation field and contact probabilities, we further propose disentangling the regression of deformation from the pose and shape of hand and face represented by mesh vertex positions into two network branches, which enhances the estimation of deformations and contacts while resulting in accurate and robust hand and face mesh recovery. Instead of directly regressing hand and face parameters, we learn an intermediate non-parametric mesh representation. We then use this representation to regress the pose and shape parameters of hand and face with a neural inverse-kinematics network. Compared with directly regressing the pose and shape parameters which learns the abstract parameters is a highly non-linear process and suffers from image-model misalignment, predicting vertex positions in Euclidean space and then applying inverse-kinematics enhances the reconstruction accuracy [53, 52, 51]. Consequently, our model achieves higher reconstruction accuracy than all previous regression and optimization-based methods. It also reaps the benefits of an animatable parametric hand and face representation that could be readily used by downstream applications.
Meanwhile, despite containing rich annotations, the existing benchmark dataset [97] collected in a studio is still limited in the diversity of hand motions, facial expressions, and appearances. Training a model only on such a dataset limits its generalization capability when applied to in-the-wild data. To achieve robust and generalizable hand-face interaction and deformation recovery, we introduce a weak-supervision training pipeline that utilizes in-the-wild images without the reliance on 3D annotations.
@@
In addition to the 2D keypoint supervision for in-the-wild images, we propose a novel depth supervision pipeline. This pipeline leverages the robust depth prior from a diffusion-based monocular depth estimation model [44], which provides essential geometric information for accurate mesh recovery and captures spatial relationships critical for contact state and deformation estimation. To improve our model’s robustness, we further employ pose priors of the hand and face by introducing hand and face parameter discriminators that learn rich hand and face motion priors from multiple datasets on hand or face separately [79, 140]. By incorporating a small set of real-world images alongside the Decaf dataset and leveraging our weak-supervision pipeline, we markedly enhance the accuracy and generalization capacity of our model.
@@ 2 Related Work
Extensive efforts have been made to recover meshes from monocular images, including human bodies [2, 71, 53, 5, 14, 116, 112, 111, 64, 43, 4, 128, 25, 59, 110, 18, 11, 40, 63], hands [93, 73, 72, 70, 76, 81, 121, 120, 54, 126], and faces [24, 26, 115, 16, 139, 7, 134, 78, 35, 8, 48, 50]. This also includes recovering the surrounding environments [12, 39, 32, 33, 135, 58, 133, 96, 69, 114] and interacting objects [119, 129, 85, 104, 30, 100, 129, 27, 86, 34, 123, 9, 10, 65, 15] while reconstructing the mesh. The acquired versatile behaviors play a crucial role in various applications, including motion generation [101, 83, 80, 28, 108, 117, 118, 61, 138, 105, 84, 17, 106], augmented reality (AR), virtual reality (VR), and human behavior analysis [130, 122, 132, 131, 29, 66]. In the following, we mainly review the related works on hand, face and full-body mesh recovery.
@@
3D Interacting Hands Recovery. Recent advancements have markedly enhanced the capture and recovery of 3D hand interactions. Early studies have achieved reconstruction of 3D hand-hand interactions utilizing a fitting framework, employing resources such as RGBD sequences [77], hand segmentation maps [74], and dense matching maps [107]. The introduction of large-scale datasets for interacting hands [73, 72] has motivated the development of regression-based approaches. Notably, these include regressing 3D interacting hand directly from monocular RGB images [93, 70, 127, 55, 141]. Additionally, research has extended to recovering interactions between hands and various objects in the environment, including rigid [6, 27, 65, 100, 20, 125, 124, 13], articulated [21], and deformable [103] objects. Following [97], our work distinguishes itself by introducing hand interactions with a deformable face, characterized by its non-uniform stiffness—a significant difference from conventional deformable models. This innovation presents unique challenges in accurately modeling interactions.
@@
3D Human Face Recovery. Research in human face recovery encompasses both optimization-based [1, 102] and regression-based [26, 94] methodologies. Beyond mere geometry reconstruction, recent approaches have evolved to incorporate training networks with the integration of differentiable renderers [24, 139, 137, 109, 11]. These methods estimate variables such as lighting, albedo, and normals to generate facial images and compare them with the monocular input. However, a significant limitation in much of the existing literature is the neglect of the face’s deformable nature and hand-face interactions. Decaf [97] represents a pivotal development in this area, attempting to model the complex mimicry of musculature and the underlying skull anatomy through optimization techniques. In contrast, our work introduces a regression-based, end-to-end method for efficient problem-solving, setting a new benchmark in the field.
@@
3D Full-Body Recovery. The task of monocular human pose and shape estimation involves reconstructing a 3D human body from a single-color image. Optimization-based approaches [2, 82, 95, 90] employ the SMPL model [67], fitting it to 2D keypoints detected within the image. Conversely, regression-based methods [53, 49, 45, 43, 23, 22, 62, 5, 25] leverage deep neural networks to directly infer the pose and shape parameters of the SMPL model. Hybrid methods [46] integrate both optimization and regression techniques, enhancing 3D model supervision. Distinct from these approaches, we follow parametric methods [53, 5, 43, 2] due to its flexibility for animation purposes. Unlike most research in this domain, which primarily concentrates on the main body with only rough estimations of hands and face, our methodology uniquely accounts for detailed interactions between these components.
@@ 3 Method
Figure 2: Overview of the proposed DICE framework. The input image is first fed to a CNN to extract a feature map, which is then passed to the Transformer-based encoders for mesh and interaction, i.e., MeshNet and InteractionNet. The MeshNet extracts hand and face mesh features, which is then used by the Inverse Kinematics models (IKNets) to predict pose and shape parameters that drive FLAME [57] and MANO [92] models. The InteractionNet predicts per-vertex hand-face contact probabilities and face deform fields from the feature map, the latter is applied to the face mesh output by the FLAME model. To improve the generalization capability, we introduce a Weakly-Supervised Training Scheme using off-the-shelf 2D keypoint detection models [68, 3] and depth estimation models [44] to provide depth supervision on keypoints. In addition, we use head and hand discriminators to constrain the distribution of parameters regressed by IKNets.
Problem Formulation. Following Decaf [97], we adopt the FLAME [57] and MANO [92] parametric models for hand and face. Given a single RGB image $\mathbf{I}\in\mathbb{R}^{224\times 224\times 3}$, the objective of this task is to reconstruct the vertices of hand mesh $\mathbf{V}_{H}\in\mathbb{R}^{778\times 3}$ and face mesh $\mathbf{V}_{F}\in\mathbb{R}^{5023\times 3}$, along with capturing the face deformation vectors $\mathbf{D}\in\mathbb{R}^{5023\times 3}$ caused by hand-face interaction and its non-rigid nature, and per-vertex contact probabilities of hand $\mathbf{C}_{H}\in\mathbb{R}^{778}$ and face ${\mathbf{C}_{F}}\in\mathbb{R}^{5023}$.
@@ 3.1 Transformer-based Hand-Face Interaction Recovery
Our model incorporates a two-branch Transformer architecture and integrates inverse-kinematic models, specifically MeshNet, InteractionNet, and IKNets. A differentiable renderer [89] is used to compute depth maps from the predicted mesh for depth supervision, and the hand and face discriminators are used as priors for constraining the hand and face poses; See Fig. 2 for an overview.
@@
Given a monocular RGB image $\mathbf{I}$, we use a pretrained HRNet-W64 [99] backbone to extract a feature map $\bm{X}_{\text{I}}\in\mathbb{R}^{H\times W\times C}$, where $H,W$ are the spatial dimension and $C$ is the channel dimension. Following [63, 64], we flatten the image feature maps and upsample the $H\times W$ feature maps to $N$ feature maps, one for each keypoint and coarse vertex of both hand and face. We then concatenate the $\mathbf{F}^{\prime}\in\mathbb{R}^{N\times C}$-dimensional feature maps with a $N\times 3$-dimensional downsampled hand and face vertices and keypoints of the mean pose, as mesh vertices and joints queries for the transformer, resulting in a feature map $\mathbf{F}\in\mathbb{R}^{N\times(C+3)}$. The mesh vertices and joints also serve as the positional encoding. To effectively model the interaction between the vertices, we mask the image feature maps corresponding to a random subset of vertices.
@@
We propose to use two separate branches, MeshNet and InteractionNet, splitting the regression of mesh vertices and a deformation field for their semantic difference: the mesh positions are more global while the deformation vectors and contact states are relatively local. Specifically, the network is followed by two progressively downsampling transformers: MeshNet, which takes the feature map $\mathbf{F}$ as input and regresses the rough vertex positions of hand $\mathbf{V}_{H}^{\prime}$ and face, $\mathbf{V}_{F}^{\prime}$; and InteractionNet, which first downsamples the feature map $\mathbf{F}$, then uses it to predict the 3D deformation field $\mathbf{D}$ at each face vertex along with the contact labels for each hand and face vertices, $\mathbf{C}_{H}$ and ${\mathbf{C}_{F}}$. Note the contacts and deformations are regressed in the same encoder to model their close relationship: the contacts cause the deformations. We validate our design in Sec. 4.4.
@@
Next, instead of directly regressing the hand and face vertices, we regress the pose and shape of the parametric hand and face models, making the output readily animatable for downstream applications. This is achieved by a neural inverse kinematics model, named IKNet, similar to Kolotouros et al. [47]. Our IKNet takes roughly estimated hand and face mesh vertices $\mathbf{V}_{H}^{\prime}$ and $\mathbf{V}_{F}^{\prime}$ as inputs and predict hand and face pose, shape and expression parameters $(\theta_{\text{h}},\beta_{\text{h}})$, $(\theta_{\text{f-pose}},\beta_{\text{f}},\theta_{\text{f-exp}})$, along with the position and orientation for hand [57] and face [92], respectively. We use the predicted parameters to first obtain the hand mesh and undeformed face mesh $\mathbf{V}_{H}$, $\mathbf{V}_{F}^{*}$. Then, we apply the deformation $\mathbf{D}$ predicted by the InteractionNet on $\mathbf{V}_{F}^{*}$ to get the final deformed face $\mathbf{V}_{F}$. Regressing parameters offers several advantages: first, it enables readily animatable meshes; second, compared to non-parametric regression methods, where meshes typically contain artifacts such as spikes [63, 11, 64], the mesh quality is significantly improved; third, the compact parameter space facilitates a more effective discriminator, which will be discussed in the following section.

### v2

@@ Abstract
Reconstructing 3D hand-face interactions with deformations from a single image is a challenging yet crucial task with broad applications in AR, VR, and gaming. The challenges stem from self-occlusions during single-view hand-face interactions, diverse spatial relationships between hands and face, complex deformations, and the ambiguity of the single-view setting. The previous state-of-the-art, Decaf, employs a global fitting optimization guided by contact and deformation estimation networks trained on studio-collected data with 3D annotations. However, Decaf suffers from a time-consuming optimization process and limited generalization capability due to its reliance on 3D annotations of hand-face interaction data. To address these issues, we present DICE, the first end-to-end method for Deformation-aware hand-face Interaction reCovEry from a single image. DICE estimates the poses of hands and faces, contacts, and deformations simultaneously using a Transformer-based architecture. It features disentangling the regression of local deformation fields and global mesh vertex locations into two network branches, enhancing deformation and contact estimation for precise and robust hand-face mesh recovery. To improve generalizability, we propose a weakly-supervised training approach that augments the training set using in-the-wild images without 3D ground-truth annotations, employing the depths of 2D keypoints estimated by off-the-shelf models and adversarial priors of poses for supervision. Our experiments demonstrate that DICE achieves state-of-the-art performance on a standard benchmark and in-the-wild data in terms of accuracy and physical plausibility. Additionally, our method operates at an interactive rate (20 fps) on an Nvidia 4090 GPU, whereas Decaf requires more than 15 seconds for a single image. The code will be available at: https://github.com/Qingxuan-Wu/DICE.
@@ 1 Introduction
Hand-face interaction is a common behavior observed up to 800 times per day across all ages and genders (Spille et al., 2021). Therefore, faithfully recovering hand-face interactions with plausible deformations is an important task given its wide applications in AR/VR (Pumarola et al., 2018; Hu et al., 2017; Wei et al., 2019), character animation (Qin et al., 2023; Zhao et al., 2024), and human behavior analysis (Liu et al., 2022; Guo et al., 2023; Mueller et al., 2019). Given the speed requirement of downstream applications like AR/VR, fast and accurate 3D reconstruction of hand-face interactions is highly desirable. However, several challenges make monocular hand-face deformation and interaction recovery particularly challenging: 1) self-occlusions involved in hand-face interaction, 2) the diversity of hand and face poses, contacts, and deformations, and 3) ambiguity in the single-view setting. Most existing methods (Rempe et al., 2020; Muller et al., 2021) only reconstruct hand (Romero et al., 2022) and face (Li et al., 2017) meshes, unified as a whole-body model (Loper et al., 2023; Pavlakos et al., 2019), without capturing contacts and deformations. A seminal advance, Decaf (Shimada et al., 2023), recovers hand-face interactions while accounting for both deformations and contacts. However, Decaf requires time-consuming optimization, which takes more than 15 seconds per image, rendering it unsuitable for interactive applications. Moreover, Decaf’s iterative fitting process depends heavily on accurate initial estimates of hand and face keypoints, as well as contact points on their surfaces, which could fail when significant occlusion is present in the image (See Fig. 8). Additionally, Decaf cannot scale up their training to fruitful hand-face interaction data in the wild, as they require 3D ground-truth annotations, such as contact labels and deformations that are not available from the in-the-wild data.
@@
Figure 1: Our method is the first end-to-end approach that captures hand-face interaction and deformation from a monocular image. Results are from (a) Decaf’s validation dataset, (b) in-the-wild images, and (c) VR use cases.
@@
To tackle the issues above, we present DICE, the first end-to-end approach for Deformation-aware hand-face Interaction reCovEry from a monocular image. Our approach features three key designs: 1) Our Transformer-based model leverages the attention mechanism to capture the relationships between the hand and face. 2) Motivated by the global nature of pose and shape, as well as the local nature of the deformation field and contact probabilities–their invariance to global transformations of the hand and face–we propose disentangling the regression of global geometry and local interaction into two network branches. We evaluate this approach to enhance the estimation of deformations and contacts while ensuring accurate and robust recovery of hand and face meshes. 3) Instead of directly regressing the hand and face parameters, we learn an intermediate non-parametric mesh representation. This representation is used to regress the pose and shape parameters of the hand and face using a neural inverse-kinematics network. Compared to directly regressing pose and shape parameters, which learns abstract parameters in a highly non-linear space, predicting vertex positions in Euclidean space and then applying inverse-kinematics improves the reconstruction accuracy (Li et al., 2021; 2023c; 2023b). Combining all these contributions, our model achieves higher reconstruction accuracy than all previous regression- (Feng et al., 2021a; Li et al., 2017; Lin et al., 2021a) and optimization-based (Shimada et al., 2023; Lugaresi et al., 2019; Li et al., 2017) methods. Additionally, by utilizing the neural inverse-kinematics network, our approach benefits from an animatable parametric representation of the hand and face, which can be readily utilized in downstream applications.
@@
Despite containing rich 3D annotations, the existing benchmark dataset (Shimada et al., 2023) collected in a studio is still limited in the diversity of hand motions, facial expressions, and appearances. Training only on such a dataset limits the model’s ability to generalize to in-the-wild scenarios. To achieve robust and generalizable hand-face interaction and deformation recovery, we introduce a weak-supervision training pipeline that utilizes in-the-wild images without the reliance on 3D annotations. To achieve this, our key insight is to leverage additional prior knowledge, such as depth supervision alongside 2D keypoint supervision, compensating for the absence of ground truth contact and deformation annotations. We leverage the robust depth prior provided by a diffusion-based monocular depth estimation model (Ke et al., 2024), which provides essential geometric information for accurate mesh recovery and captures spatial relationships critical for contact state and deformation estimation. As the task becomes highly ill-posed for in-the-wild images, we further employ pose priors of the hand and face by introducing hand and face parameter discriminators that learn rich hand and face motion priors from additional datasets on hand or face separately (Pan et al., 2023a; Zimmermann et al., 2019). By incorporating a small set of real-world images alongside the Decaf dataset and leveraging our weak-supervision pipeline, we markedly enhance the accuracy and generalization capacity of our model.
@@ 2 Related Work
Extensive efforts have been made to recover meshes from monocular images, including human bodies (Bogo et al., 2016; Moon & Lee, 2020; Li et al., 2021; Cai et al., 2024; Contributors, ; Xie et al., 2022; Wang & Daniilidis, 2023; Wang et al., 2023b; Lin et al., 2021b; Kanazawa et al., 2018; Cai et al., 2022; Zhang et al., 2021b; Feng et al., 2023; Li et al., 2022c; Wang et al., 2023a; Dou et al., 2023b; Cho et al., 2022; Huang et al., 2022b; Lin et al., 2021a), hands (Rong et al., 2021; Moon et al., 2020; 2024; Moon, 2023; Oh et al., 2023; Park et al., 2022; Yang et al., 2021; 2022b; Li et al., 2023d; Yu et al., 2023), and faces (Feng et al., 2021b; 2018; Wood et al., 2022; Daněček et al., 2022; Zielonka et al., 2022; Chai et al., 2023; Zhang et al., 2023c; Otto et al., 2023; He et al., 2023; Chatziagapi & Samaras, 2023; Kumar et al., 2023; Li et al., 2023a). This also includes recovering the surrounding environments (Clever et al., 2022; Huang et al., 2022a; Hassan et al., 2019; 2021; Zhang et al., 2020b; Li et al., 2022b; Zhang et al., 2021c; Shimada et al., 2022; Luo et al., 2022; Weng & Yeung, 2021) and interacting objects (Yang et al., 2022a; Zhang et al., 2020a; Pham et al., 2017; Tsoli & Argyros, 2018; Hampali et al., 2020; Tekin et al., 2019; Zhang et al., 2020a; Grady et al., 2021; Pokhariya et al., 2023; Hasson et al., 2019; Ye et al., 2022; Chen et al., 2023; 2021; Liu et al., 2021; Corona et al., 2020) while reconstructing the mesh. The acquired versatile behaviors play a crucial role in various applications, including motion generation (Tevet et al., 2022; Peng et al., 2022; Pan et al., 2023b; Guo et al., 2022; Wang et al., 2022a; Xu et al., 2023; 2024; Lin et al., 2024; Zhou et al., 2023; Wan et al., 2023a; Peng et al., 2021; Dou et al., 2023a; Wan et al., 2023b), augmented reality (AR), virtual reality (VR), and human behavior analysis (Zhang et al., 2023a; Yang et al., 2024; Zhang et al., 2024; 2023b; Guo et al., 2023; Liu et al., 2022). In the following, we mainly review the related works on hand, face and full-body mesh recovery.
@@
3D Interacting Hands Recovery. Recent advancements have markedly enhanced the capture and recovery of 3D hand interactions. Early studies have achieved reconstruction of 3D hand-hand interactions utilizing a fitting framework, employing resources such as RGBD sequences (Oikonomidis et al., 2012), hand segmentation maps (Mueller et al., 2019), and dense matching maps (Wang et al., 2020). The introduction of large-scale datasets for interacting hands (Moon et al., 2020; 2024) has motivated the development of regression-based approaches. Notably, these include regressing 3D interacting hand directly from monocular RGB images (Rong et al., 2021; Moon, 2023; Zhang et al., 2021a; Li et al., 2022a; Zuo et al., 2023). Additionally, research has extended to recovering interactions between hands and various objects in the environment, including rigid (Cao et al., 2021; Grady et al., 2021; Liu et al., 2021; Tekin et al., 2019; Fan et al., 2024; Ye et al., 2023b; a), articulated (Fan et al., 2023), and deformable (Tretschk et al., 2023) objects. Following Shimada et al. (2023), our work distinguishes itself by introducing hand interactions with a deformable face, characterized by its non-uniform stiffness—a significant difference from conventional deformable models. This innovation presents unique challenges in accurately modeling interactions.
@@
3D Human Face Recovery. Research in human face recovery encompasses both optimization-based (Aldrian & Smith, 2012; Thies et al., 2016) and regression-based (Feng et al., 2018; Sanyal et al., 2019) methodologies. Beyond mere geometry reconstruction, recent approaches have evolved to incorporate training networks with the integration of differentiable renderers (Feng et al., 2021b; Zielonka et al., 2022; Zheng et al., 2022; Wang et al., 2022b; Cho et al., 2022). These methods estimate variables such as lighting, albedo, and normals to generate facial images and compare them with the monocular input. However, a significant limitation in much of the existing literature is the neglect of the face’s deformable nature and hand-face interactions. Decaf (Shimada et al., 2023) represents a pivotal development in this area, attempting to model the complex mimicry of musculature and the underlying skull anatomy through optimization techniques. In contrast, our work introduces a regression-based, end-to-end method for efficient problem-solving, setting a new benchmark in the field.
@@
3D Full-Body Recovery. The task of monocular human pose and shape estimation involves reconstructing a 3D human body from a single image. Optimization-based approaches (Bogo et al., 2016; Pavlakos et al., 2019; Shi et al., 2023; Rempe et al., 2021) employ the SMPL model (Loper et al., 2023), fitting it to 2D keypoints detected within the image. Conversely, regression-based methods (Li et al., 2021; Lassner et al., 2017; Kocabas et al., 2021; Kanazawa et al., 2018; Feng et al., 2021a; Fang et al., 2021; Lin et al., 2023; Cai et al., 2024; Feng et al., 2023) leverage deep neural networks to directly infer the pose and shape parameters of the SMPL model. Hybrid methods (Kolotouros et al., 2019a) integrate both optimization and regression techniques, enhancing 3D model supervision. Distinct from these approaches, we follow parametric methods (Li et al., 2021; Cai et al., 2024; Kanazawa et al., 2018; Bogo et al., 2016) due to its flexibility for animation purposes. Unlike most research in this domain, which primarily concentrates on the main body with only rough estimations of hands and face, our methodology uniquely accounts for detailed interactions between these components.
@@ 3 Method
Figure 2: Overview of the proposed DICE framework. The input image is first fed to a CNN to extract a feature map, which is then passed to the Transformer-based encoders for mesh and interaction, i.e., MeshNet and InteractionNet. MeshNet extracts hand and face mesh features, which are then used by the Inverse Kinematics models (IKNets) to predict pose and shape parameters that drive FLAME (Li et al., 2017) and MANO (Romero et al., 2022) models. InteractionNet predicts per-vertex hand-face contact probabilities and face deformation fields from the feature map, where the latter is applied to the face mesh output by the FLAME model. To improve the generalization capability, we introduce a weakly-supervised training scheme using off-the-shelf 2D keypoint detection models (Lugaresi et al., 2019; Bulat & Tzimiropoulos, 2017) and depth estimation models (Ke et al., 2024) to provide depth supervision on keypoints. In addition, we use face and hand discriminators to constrain the distribution of parameters regressed by IKNets.
Problem Formulation. Following Decaf (Shimada et al., 2023), we adopt the FLAME (Li et al., 2017) and MANO (Romero et al., 2022) parametric models for hand and face. Given a single RGB image $\mathbf{I}\in\mathbb{R}^{224\times 224\times 3}$, the objective of this task is to reconstruct the vertices of a hand mesh $\mathbf{V}_{H}\in\mathbb{R}^{778\times 3}$ and a face mesh $\mathbf{V}_{F}\in\mathbb{R}^{5023\times 3}$, along with capturing the face deformation vectors $\mathbf{D}\in\mathbb{R}^{5023\times 3}$ resulting from hand-face interaction and its non-rigid nature. Additionally, we estimate per-vertex contact probabilities of hand $\mathbf{C}_{H}\in\mathbb{R}^{778}$ and face ${\mathbf{C}_{F}}\in\mathbb{R}^{5023}$.
@@ 3.1 Transformer-based Hand-Face Interaction Recovery
Our model incorporates a two-branch Transformer architecture and integrates inverse-kinematic models, specifically, MeshNet, InteractionNet, and IKNets. A differentiable renderer (Ravi et al., 2020) is used to compute depth maps from the predicted mesh for depth supervision, while the hand and face discriminators are used as priors for constraining the hand and face poses; See Fig. 2 for an overview.
@@
Given a monocular RGB image $\mathbf{I}$, we use a pretrained HRNet-W64 (Sun et al., 2019) backbone to extract a feature map $\bm{X}_{\text{I}}\in\mathbb{R}^{H\times W\times C}$. Following Lin et al. (2021a; b), we flatten the image feature maps and upsample the $H\times W$ feature maps to $N$ feature maps, corresponding to each keypoint and downsampled vertex of both hand and face. The feature maps $\mathbf{F}^{\prime}\in\mathbb{R}^{N\times C}$ are then concatenated with the downsampled hand and face vertex and keypoint coordinates of dimension $N\times 3$, with the pose set to the mean pose, serving as positional encodings. This results in the final feature map $\mathbf{F}\in\mathbb{R}^{N\times(C+3)}$. To model the vertex-vertex interactions, we mask the feature maps $\mathbf{F}$ for a randomly selected subset of vertices.
@@
Once the feature map $\mathbf{F}$ is obtained, it is fed into MeshNet and InteractionNet, which handle the regression of mesh vertices and the deformation field separately. This decomposition is motivated by their semantic differences: mesh contains global features, whereas deformation vectors and contact states are localized features, i.e., invariant to the global transformations of the hand and face. Thus, MeshNet takes the feature map $\mathbf{F}$ as input and regresses the unrefined vertex positions of hand $\mathbf{V}_{H}^{\prime}$ and face $\mathbf{V}_{F}^{\prime}$. InteractionNet, on the other hand, predicts the 3D deformation field $\mathbf{D}$ for each face vertex, along with the contact labels for each hand and face vertex, $\mathbf{C}_{H}$ and $\mathbf{C}_{F}$, respectively. Note the contacts and deformations are regressed within the same encoder to model their causal relationship: the contacts cause the deformations. We validate our design in Sec. 4.4.
@@
Next, instead of directly using the unrefined hand and face vertices $\mathbf{V}_{H}^{\prime}$ and $\mathbf{V}_{F}^{\prime}$, our method takes these vertices as input to regress the pose and shape of their respective parametric models  (Li et al., 2017; Romero et al., 2022). This is achieved by a neural inverse kinematics model, named IKNet, following Kolotouros et al. (2019b). The IKNet takes the unrefined hand and face vertices $\mathbf{V}_{H}^{\prime}$ and $\mathbf{V}_{F}^{\prime}$ as inputs and predicts their pose, shape, and expression parameters $(\bm{\theta}_{\text{h}},\bm{\beta}_{\text{h}})$ for hand, $(\bm{\theta}_{\text{f-pose}},\bm{\beta}_{\text{f}},\bm{\theta}_{\text{f-exp}})$ for face, along with the root position and orientation for hand $(\bm{t}_{h},\bm{r}_{h})$ and face $(\bm{t}_{f},\bm{r}_{f})$, respectively. Afterward, we use the predicted parameters to first obtain the hand vertices $\mathbf{V}_{H}$ and undeformed face vertices $\mathbf{V}_{F}^{*}$. Then, we apply the deformation $\mathbf{D}$ predicted by the InteractionNet on $\mathbf{V}_{F}^{*}$ to get the final deformed face $\mathbf{V}_{F}$. Utilizing parametric forward-kinematics and neural inverse-kinematics models offer several advantages: first, it enables readily animatable meshes for downstream applications; second, compared to non-parametric regression methods, where meshes typically contain artifacts such as spikes (Lin et al., 2021a; Cho et al., 2022; Lin et al., 2021b), this approach significantly improves mesh quality; third, the compact parameter space allows for a more effective discriminator, which will be discussed in the following section.